OpenStudy (anonymous):

4. Bill is able to save $35/week after working part-time and paying his expenses. These two formulas show his weekly savings: A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 B. f(n) = 35n 1. Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. 2. Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (anonymous):

use this rule A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 make a table n | f(n) ----------- 1 | 35 2 | ? you need to find f(2) the rule says, for n>1 (and here n is 2, so we use it) f(n) = f(1)+ f(n-1) replace n with 2: f(2) = f(1) + f(2-1) or f(2) = f(1)+f(1) from the table, we see f(1) is 35, so you get f(2) = 35+35 f(2) = 70

OpenStudy (anonymous):

now fill in n=3, and find f(3) can you do that ?

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

got it thanks!

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

Given the sequence of numbers: 5, 6, 8, 11, 15, 20, 26, 33, 41,… Explain whether or not this sequence can be considered a function.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2. Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations.